. During the Great Recession, U.S investors suffered a huge loss in business confidence once the stock market crashed in 2008. Consider the model below: Y = 1650; C = 200 + 0.6(Y – T); G = 250; I = 400 – 2000r; T= 150 i. Find the equilibrium interest rate (r), investment (I), and consumption (C), public savings, private savings, and national savings.
Answer #1

At equilibrium Y = c ta ti 1650 = 2009 to.6(1650 - 150) +250t 400 - 2000s 100 = 2000 & r= 0.05 C = 200 +0.6 (1650 - 150) = 11